OpenStudy (anonymous):

What is the axis of symmetry? graph of y = 2(x – 3)2 – 7

hero (hero):

whatevs

OpenStudy (anonymous):

hey

hero (hero):

Hi

hero (hero):

There's probably an easier way to do this, but I usually just multiply everything all out first

OpenStudy (anonymous):

ok

hero (hero):

y = 2(x – 3)^2 – 7 y = 2(x^2-6x+9) - 7 y = 2x^2-12x + 18 - 7 y = 2x^2-12x+11 Now, in order to find the axis of symmetry, we first have to find the vertex... x = -b/(2a)

OpenStudy (anonymous):

12/4 = 3

OpenStudy (anonymous):

so 3 is the vertex?

hero (hero):

-(-12)/2(2) = 12/4 = 3 x = 3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the axis of symmetry should be 3 right?

hero (hero):

yes

OpenStudy (anonymous):

thanks